maria company has 6000 shares of $.50 par value common stock and 900 shares of $3 cumulative preferred stock outstanding. The dividends on the preferred stock were not declared during the previous 3 years. THis year the company is declaring dividends of $16,800. The common stockholders receive how much in dividends per share?
Answer #1

Number of preferred shares = 900

Dividend per preferred share = $3

Annual preferred dividend = Number of preferred shares x Dividend per preferred share

= 900 x 3

= $2,700

Preferred dividend have not been declared for the last 3 years.

Arrears of preferred dividends = Annual preferred dividend x 3

= 2,700 x 3

= $8,100

Preferred dividends to be paid this year = Arrears of preferred dividends+ Annual preferred dividend

= 8,100+2,700

= $10,800

Total dividends declared this year = $16,800

Dividend received by common stockholders = Total dividends declared this year- Preferred dividends to be paid this year

= 16,800-10,800

= $6,000

Dividend per share of common stock = Dividend received by common stockholders/ Number of common shares

= 6,000/6,000

= $1

The common stockholders receive $1 dividends per share.

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!